# A Deep Learning Method for Simultaneous Denoising and Missing Wedge Reconstruction in Cryogenic Electron Tomography

### v2

@@ 2.2 Experiments
We now study the performance of DeepDeWedge on real and synthetic data. We compare DeepDeWedge to CryoCARE, IsoNet, and a two-step approach of fitting IsoNet to tomograms denoised with CryoCARE. The two-step approach, which we call CryoCARE + IsoNet, is considered a state-of-the-art pipeline for denoising and missing wedge reconstruction. We begin with qualitative experiments on three real-world datasets and then quantitatively evaluate differences between the methods by experiments on synthetic data.
We find that DeepDeWedge performs similarly, and in some cases better than CryoCARE, IsoNet and CryoCARE + IsoNet, in terms of both denoising and missing wedge reconstruction.
2.2.1 Experiments on Purified S. Cerevisiae 80S Ribosomes (EMPIAR-10045)
Figure 3: Slices through 3D reconstructions of a tomogram containing purified S. cerevisiae 80S ribosomoes (EMPIAR-10045, Tomogram 5).
The first dataset we consider is EMPIAR-10045, the tutorial dataset for sub-tomogram averaging in Relion [BS16]. We chose this dataset because it is commonly used to test new algorithms, for example to evaluate IsoNet [Liu+22].
The EMPIAR-10045 dataset contains 7 tilt series collected from samples of purified S. Cerevisiae 80S Ribosomes. All tilt series are aligned and consist of 41 projections collected at tilt angles from $-60^{\circ}$ to $+60^{\circ}$ with a $3^{\circ}$ increment. The original pixel size of the projections is 2.17 Å. After FBP reconstruction, we binned all tomograms 6 times, resulting in a cubic voxel size of $13.02$. Finally, we applied IsoNet’s Wiener-filter-based CTF correction to all tomograms. We fitted all models using sub-tomograms with shape $80\times 80\times 80$.
@@
For IsoNet, we used the authors’ original implementation (https://github.com/IsoNet-cryoET/IsoNet), and for CryoCARE, we used our own re-implementation. As model architecture, we used a U-Net [RFB15] with 64 channels and 3 downsampling layers for all three methods. The network has $27.3$ million trainable parameters and is also the default model architecture of IsoNet.
@@
By default, IsoNet uses regularization with dropout with probability $0.3$, which we also kept for CryoCARE to prevent overfitting. For DeepDeWedge, we found that dropout is unnecessary and increases the fitting time, so we did not use dropout for our method. We fitted all models using the Adam optimizer [KB15] with a constant learning rate of $4\cdot 10^{-4}$. For more details on the experiment, please refer to Appendix D.1.
@@
Results: Figure 3 shows tomograms refined with IsoNet, CryoCARE + IsoNet and DeepDeWedge using the even/odd tilt series splitting approach. Note that while IsoNet’s built-in noisier2noise-like denoiser removes some of the noise contained in the FBP reconstruction, its performance is considerably worse than that of noise2noise-based CryoCARE. This can be seen by comparing to the result of applying IsoNet with disabled denoiser to a tomogram denoised with CryoCARE (CryoCARE + IsoNet). DeepDeWedge produces a denoised and missing-wedge-corrected tomogram similar to the Cryocare + IsoNet combination. The main difference between these reconstructions is that the DeepDeWedge-refined tomogram has a smoother background and contains fewer high-frequency components.
@@
Regarding missing wedge correction, we find the performance of DeepDeWedge and IsoNet to be similar. In slices parallel to the x-z-plane, where the effects of the missing wedge on the FBP reconstruction are most prominent, both IsoNet and DeepDeWedge reduce artifacts and correct artificial elongations of the ribosomes.
2.2.2 Experiments on Flagella of Chlamydomonas Reinhardtii
Next, we evaluate DeepDeWedge on a second real-world dataset of tomograms of the flagella of Chlamydomonas reinhardtii, which is the tutorial dataset for CryoCARE. Since we observed above that IsoNet performs better when applying it to CryoCARE-denoised tomograms, we compare only to CryoCARE + IsoNet. In addition, we investigate the impact of splitting the tilt-series into even and odd projections versus using the frame-based split for DeepDeWedge and CryoCARE + IsoNet.
The dataset is a single tilt series collected from the flagella of Chlamydomonas reinhardtii. The projections were collected at angles from $-65^{\circ}$ to $+65^{\circ}$ with $2^{\circ}$ increment. Each projection was acquired using dose fractionation with $10$ frames per tilt angle. Each frame has pixel size $2.36$. Again, a binning factor of 6 was used for reconstruction, resulting in a voxel size of $14.16$. Finally, we extracted sub-tomograms of size $96\times 96\times 96$ from the FBP reconstruction to fit all models. For more details on the experiment, please refer to Appendix D.2.
@@ 2.2.2 Experiments on Flagella of Chlamydomonas Reinhardtii
Figure 4: Slices through 3D reconstructions of the flagella of Chlamydomonas reinhardtii when using different reconstruction methods.
@@
The reconstructions obtained with DeepDeWege and CryoCare + IsoNet are shown in Figure 4. We find that when using the even/odd-based split, the Cryocare + IsoNet combination produces a more crisp reconstruction compared to DeepDeWedge (see zoomed in region). This may be due to the fact that the model inputs and targets of IsoNet stem from denoised FBP reconstruction of the full tilt series in which the information is more densely sampled than in the sub-tomograms of the even and odd FBP to the reconstructions used for fitting DeepDeWedge. When using the frame-based splitting method, in which DeepDeWedge also operates on the more densely sampled FBP reconstruction, the denoising performance is on par with that of Cryocare + IsoNet.
Both IsoNet and DeepDeWedge fix the missing-wedge-caused distortions of the microtubules exhibited by the FBP reconstruction, as seen in the x-z slices. There, the DeepDeWedge reconstruction has higher contrast and is less noisy than that produced by Cryocare + IsoNet for both the even/odd and the frame-based splitting approach. Moreover, DeepDeWedge reconstructed more of the flagellas’ outer parts.
2.2.3 Experiments on the Ciliary Transit Zone of Chlamydomonas Reinhardtii (EMPIAR-11078)
@@
Figure 5: Slices through 3D reconstructions of tomograms showing the ciliary transit zone of Chlamydomonas reinhardtii. Each row shows one of the three tomograms we used for model fitting. The row labels follow the naming convention of EMPIAR-11078.
@@
Finally, we apply DeepDeWedge to an in situ dataset. We chose EMPIAR-11078 [Hoe+22], which contains tilt series of the ciliary transit zone of Chlamydomonas reinhardtii. The crowded cellular environment together with the low contrast and signal-to-noise ratio of the tilt series make EMPIAR-11078 significantly more challenging for denoising and missing wedge reconstruction than two datasets from our previous experiments.
Again, we compare DeepDeWedge to CryoCARE + IsoNet. For our comparison, we chose Tomograms 2, 3, and 8, since their tilt series were approximately collected at the angular range from $-60^{\circ}$ to $+60^{\circ}$ (increment: $2^{\circ}$), for which IsoNet was developed and optimized. We fitted DeepDeWedge and the CryoCARE denoiser for CryoCARE + IsoNet using the frame-based split of the tilt-series data. For that, we reconstructed the volumes from the raw movie frames using IMOD. In the process, we binned the tilt series 6 times, resulting in a pixel size of $20.52$. From all FBP reconstructions, we extracted sub-tomograms of size $96\times 96\times 96$ to fit all models. For more details on the experiment, please refer to Appendix D.3.
Results: Looking at Figure 5, we find the denoising performance of DeepDeWedge to be similar to that of CryoCARE + IsoNet. In the x-y-plane, the DeepDeWedge reconstructions are smoother, especially in empty or background regions.
@@
In the x-z- and z-y-planes, the DeepDeWedge reconstructions are sharper and less noisy than those produced with Cryocare + IsoNet. Especially in the x-z-plane, where the effects of the missing wedge are strongest, DeepDeWedge produces higher contrast than Cryocare + IsoNet, and removes more of the artefacts. Remarkably, as can be seen in the zoomed-in regions in the second row of Figure 5, both IsoNet and DeepDeWedge reconstruct parts of the sample which are barely present in the FBP reconstruction since they are perpendicular to the electron beam direction, which means that a large portion of their Fourier components are masked out by the missing wedge.
2.2.4 Experiments on Synthetic Data
On the real-world data, we observed that DeepDeWedge performs similar, and in some cases better than CryoCARE + IsoNet in terms of both denoising and missing wedge reconstruction. In this section, we compare DeepDeWedge to IsoNet, CryoCARE, and the Cryocare + IsoNet combination on synthetic data to quantify our qualitative findings on real data. We find that the denoising performance of DeepDeWedge is on par with that of CryoCARE in terms of qualitative denoising metrics. Regarding missing wedge reconstruction, we find that our method slightly outperforms IsoNet.
@@ 2.2.4 Experiments on Synthetic Data
We used a dataset by [Gub+20, ] containing 10 noiseless synthetic ground truth volumes with a spatial resolution of 10 Å${}^{3}$/voxel and a size of $179\times 512\times 512$ voxels. All volumes contain typical objects found in cryo-ET samples, such as proteins (up to 1,500 uniformly rotated samples from a set of 13 structures), membranes, and gold fiducial markers. We selected the first three tomograms and used the Python library tomosipo [Hen+21] to compute clean projections of size $512\times 512$ in the angular range ±$\pm 60^{\circ}$ with $2^{\circ}$ increment. From these clean tilt series, we generated datasets with different noise levels by adding pixel-wise independent Gaussian noise to the projections. We simulated three datasets with tilt series SNR $1/2$, $1/4$, and $1/6$, respectively. We used 150 sub-tomograms of size $96\times 96\times 96$ to fit all models. For more details on the experiment, please refer to Appendix D.4.
@@
Figure 6: Comparison of our DeepDeWedge (with even/odd split) to the IsoNet, CryoCARE and the combination of IsoNet and CryoCARE for increasing noise levels on the tilt series images. We the y-axes such that higher is better for all metrics.
@@
Figure 7: Slices through tomograms reconstructed with Cryocare + IsoNet and DeepDeWedge (with even/odd split) on a tomogram from the synthetic SHREC20201 dataset. The clean ground truth is the same for all noise levels.
@@
Evaluation Metrics: To measure the overall quality of a tomogram $\hat{\mathbf{v}}$ obtained with any of the three methods, we calculated the normalized correlation coefficient $\text{CC}(\hat{\mathbf{v}},\mathbf{v}^{*})$ between the reconstruction $\hat{\mathbf{v}}$ and the corresponding ground-truth $\mathbf{v}^{*}$, which is defined as
@@
Results: Figure 6 shows the metrics for decreasing SNR of the tilt series. All metrics suggest that DeepDeWedge yields higher-quality reconstructions than IsoNet, CryoCARE, and CryoCARE + IsoNet.
CryoCARE achieves a lower correlation coefficient than IsoNet in the high-SNR regime, while the order is reversed for low SNR. A likely explanation is that for lower noise levels, the correlation coefficient is more sensitive to the missing wedge artifacts in the reconstructions. CryoCARE does not perform missing wedge reconstruction, so it has a lower correlation coefficient than IsoNet for higher SNR. For lower SNR, the correlation coefficient is dominated by the noise. Regarding denoising, we and others [Mal+23] have observed that CryoCARE has better performance than IsoNet, which is confirmed here by looking at the correlation coefficient outside the missing wedge. As expected, CryoCARE + IsoNet combines the strengths of both methods. Compared to this combination, the overall quality of DeepDeWedge reconstructions is on par or better, depending on the noise level.
@@
The perceived quality of the reconstructions shown in Figure 7 agrees with the metric-based comparison.
@@ 3 Discussion
In this paper, we have proposed DeepDeWedge, which is a new deep-learning-based method for denoising and missing wedge reconstruction in cryo-ET. The missing wedge-filling ability of DeepDeWedge is on par with that of its closest competitor, IsoNet. At the same time, DeepDeWedge achieves significantly improved denoising performance that matches state-of-the-art noise2noise-like denoising methods such as CryoCARE, which do not perform missing wedge reconstruction. Compared to the combination of CryoCARE and IsoNet, which is the state-of-the-art pipeline for denoising and missing wedge reconstruction, DeepDeWedge achieves similar or even better performance. At the same time, DeepDeWedge has fewer hyperparameters and is simpler as it fits only a single network for both denoising and missing wedge reconstruction simultaneously rather than sequentially fitting two separate networks for each task.
@@
The data that DeepDeWedge fills in for the missing wedge is based on prior information about the data that the network learns during model fitting by predicting data that we artificially removed. However, it is important to keep in mind that the actual information contained in the missing wedge is irreversibly lost during tilt series acquisition. Therefore, we advise users of our method to be cautious, especially regarding reconstructing objects that are mostly perpendicular to the electron beam. Many of the Fourier components that correspond to such objects are contained in the missing wedge region. Examples of objects that are almost completely masked by the missing wedge include thin membranes or elongated proteins perpendicular to the electron beam; we discuss a concrete example and how DeepDeWedge handles it in Appendix F. It remains an open question to what extent missing-wedge-filled tomograms, whether based on deep learning or classical methods, can be trusted (see also Appendix G). To estimate the trustworthiness of a DeepDeWedge reconstruction, we recommend comparing it to a reconstruction obtained with a classical, prior-free method that is strongly data-consistent, such as FBP.
@@
Based on these considerations, we see two main applications for DeepDeWedge: First, DeepDeWedge is an effective method to improve the direct interpretability of tomograms, which may enable discoveries that are prevented by high noise levels and strong missing wedge artifacts in raw FBP reconstructions. Second, tomograms reconstructed with DeepDeWedge can be used as input for downstream tasks such as segmentation or particle picking, and might improve the effectiveness of those downstream tasks. It has already been observed that denoised and/or missing-wedge-corrected tomograms can improve the performance of deep-learning-based particle pickers [Buc+19, Liu+22, Hon+23]. For subsequent steps, such as sub-tomogram averaging, one can then use the original raw tomograms that are not modified by the neural network.
@@
All tomograms reconstructed or processed using any method discussed in this paper are available from S.W. upon request.
The datasets EMPIAR-10045 and EMPIAR-11078 datasets used for the experiments discussed in Section 2.2.1 and Section 2.2.3 are publicly available through EMPIAR: https://www.ebi.ac.uk/empiar/. The tutorial dataset of CryoCARE is publicly available via GitHub https://github.com/juglab/cryoCARE_T2T, through the following download link: https://download.fht.org/jug/cryoCARE/Tomo110.zip. The synthetic dataset by [Gub+20, ] is also publicly available and can be downloaded from this website: https://dataverse.nl/dataset.xhtml?persistentId=doi:10.34894/XRTJMA.

### v3

@@ 3 Experiments
In the following sections, we compare DeepDeWedge to a re-implementation of CryoCARE, IsoNet, and a two-step approach of fitting IsoNet to tomograms denoised with CryoCARE. The two-step approach, which we call CryoCARE + IsoNet, is considered a state-of-the-art pipeline for denoising and missing wedge reconstruction.
We implemented DeepDeWedge in Python, using PyTorch [28] as a deep learning framework. For IsoNet, we used the authors’ original implementation (https://github.com/IsoNet-cryoET/IsoNet), and for CryoCARE, we used our own re-implementation, which we based on our implementation of DeepDeWedge.
@@
As model architecture, we used a U-Net [14] with 64 channels and 3 downsampling layers for all three methods. The network has $27.3$ million trainable parameters and is also the default model architecture of IsoNet.
@@
IsoNet uses regularization with dropout with probability $0.3$, which we also kept for CryoCARE to prevent overfitting. For DeepDeWedge, we found that dropout is unnecessary, so we did not use dropout for our method unless explicitly stated otherwise.
We fitted all models using the Adam optimizer [15] with a constant learning rate of $4\cdot 10^{-4}$.
3.1 Experiments on Purified S. Cerevisiae 80S Ribosomes
Figure 3: Slices through 3D reconstructions of a tomogram containing purified S. cerevisiae 80S ribosomoes (EMPIAR-10045, Tomogram 5). The red lines in each slice indicate the positions of the remaining two slices. We also show the central x-z-slice through the logarithm of the magnitude of the Fourier transform of each reconstruction.
The first dataset we consider is the commonly used EMPIAR-10045 dataset, which contains 7 tilt series collected from samples of purified S. cerevisiae 80S Ribosomes. More details on the data and the model fitting can be found in Supplementary Information 4.1.
Results.
@@
Figure 3 shows a tomogram refined with IsoNet, CryoCARE + IsoNet and DeepDeWedge using the even/odd tilt series splitting approach. Note that while IsoNet’s built-in Noisier2Noise-like denoiser removes some of the noise contained in the FBP reconstruction, its performance is considerably worse than that of Noise2Noise-based CryoCARE. This can be seen by comparing to the result of applying IsoNet with disabled denoiser to a tomogram denoised with CryoCARE. DeepDeWedge produces a denoised and missing-wedge-corrected tomogram similar to the CryoCARE + IsoNet combination. The main difference between these reconstructions is that the DeepDeWedge-refined tomogram has a smoother background and contains fewer high-frequency components.
@@
Regarding missing wedge correction, we find the performance of DeepDeWedge and IsoNet to be similar. In slices parallel to the x-z-plane, where the effects of the missing wedge on the FBP reconstruction are most prominent, both IsoNet and DeepDeWedge reduce artifacts and correct artificial elongations of the ribosomes. The central x-z-slices through the reconstructions’ Fourier transforms confirm that all methods but FBP fill in most of the missing wedge.
3.2 Experiments on Flagella of C. Reinhardtii
@@
Figure 4: Slices through 3D reconstructions of the flagella of C. reinhardtii when using different reconstruction methods. The red lines in each slice indicate the positions of the remaining two slices. We also show the central x-z-slice through the logarithm of the magnitude of the Fourier transform of each reconstruction.
Next, we evaluate DeepDeWedge on another real-world dataset of tomograms of the flagella of C. reinhardtii, which is the tutorial dataset for CryoCARE. Since we observed above that IsoNet performs better when applying it to CryoCARE-denoised tomograms, we compare only to CryoCARE + IsoNet. In addition, we investigate the impact of splitting the tilt-series into even and odd projections versus using the frame-based split for DeepDeWedge and CryoCARE + IsoNet. More details on the data and the model fitting can be found in Supplementary Information 4.2.
Results.
@@
The reconstructions obtained with all methods are shown in Figure 4. We find that when using the even/odd-based split, CryoCARE + IsoNet produces a crisper reconstruction than DeepDeWedge (see zoomed-in region). This may be because the model inputs and targets of IsoNet stem from denoised FBP reconstruction of the full tilt series in which the information is more densely sampled than in the sub-tomograms of the even and odd FBP to the reconstructions used for fitting DeepDeWedge.
@@
Figure 5: Slices through 3D reconstructions of tomograms showing the ciliary transit zone of C. reinhardtii. The red lines in each slice indicate the positions of the remaining two slices. We also show the central x-z-slice through the logarithm of the magnitude of the Fourier transform of each reconstruction. The row labels follow the naming convention of EMPIAR-11078.
When using the frame-based splitting method, in which DeepDeWedge also operates on the more densely sampled FBP reconstruction, DeepDeWedge removes more noise than CryoCARE + IsoNet and produces higher contrast, which is most noticeable in the x-z-slice. Especially in background areas, the DeepDeWedge reconstruction has fewer high-frequency components and is smoother. Therefore, the CryoCARE + IsoNet reconstruction is slightly more faithful to the FBP reconstruction but is also noisier.
The central x-z-slices through the reconstructions’ Fourier transforms’ indicate that both CryoCARE + IsoNet and DeepDeWedge fill in most of the missing wedge. Both methods fix the missing-wedge-caused distortions of the microtubules exhibited by the FBP reconstruction, as seen in the x-z slices. DeepDeWedge reconstructs more of the flagellas’ outer parts.
3.3 Experiments on the Ciliary Transit Zone of C. Reinhardtii
@@
Finally, we apply DeepDeWedge to an in situ dataset. We chose EMPIAR-11078 [29], which contains tilt series of the ciliary transit zone of C. reinhardtii. The crowded cellular environment and low contrast and signal-to-noise ratio of the tilt series make EMPIAR-11078 significantly more challenging for denoising and missing wedge reconstruction than the two datasets from our previous experiments. More details on the data and the model fitting can be found in Supplementary Information 4.3.
Results.
Slices through reconstructions of two tomograms obtained with FBP, CryoCARE + IsoNet and DeepDeWedge are shown in Figure 5. In the x-z- and z-y-planes, the DeepDeWedge reconstructions are more crisp and less noisy than those produced with CryoCARE + IsoNet. Especially in the x-z-plane, where the effects of the missing wedge are strongest, DeepDeWedge produces higher contrast than CryoCARE + IsoNet and removes more of the artifacts. Again, the CryoCARE + IsoNet reconstructions contain more high-frequency components and are closer to the FBP reconstruction, whereas the DeepDeWedge reconstructions are smoother and more denoised, especially in empty or background regions.
@@
Remarkably, as can be seen in the zoomed-in regions in the second row of Figure 5, both CryoCARE + IsoNet and DeepDeWedge reconstruct parts of the sample that are barely present in the FBP reconstruction since they are perpendicular to the electron beam direction, which means that a large portion of their Fourier components are masked out by the missing wedge.
Note that both CryoCARE + IsoNet and DeepDeWedge appear less effective at reconstructing the missing wedge compared to the two experiments presented above. This is indicated by the central x-z-slices through the Fourier transforms of the reconstructions and is likely due to the challenging, crowded nature and low SNR of the data.
3.4 Experiments on Synthetic Data
Here, we compare DeepDeWedge to IsoNet, CryoCARE, and the CryoCARE + IsoNet combination on synthetic data to quantify our qualitative findings on real data.
@@ 3.4 Experiments on Synthetic Data
We used a dataset by Gubins et al. [30] containing 10 noiseless synthetic ground truth volumes with a spatial resolution of 10 Å3/voxel and a size of $179\times 512\times 512$ voxels. All volumes contain typical objects found in cryo-ET samples, such as proteins (up to 1,500 uniformly rotated samples from a set of 13 structures), membranes, and gold fiducial markers.
For our comparison, we fitted models on the first three tomograms of the SHREC 2021 dataset. We used the Python library tomosipo [31] to compute clean projections of size $512\times 512$ in the angular range ±$\pm 60^{\circ}$ with $2^{\circ}$ increment. From these clean tilt series, we generated datasets with different noise levels by adding pixel-wise independent Gaussian noise to the projections. We simulated three datasets with tilt series SNR $1/2$, $1/4$, and $1/6$, respectively. More details on the data and the model fitting can be found in Supplementary Information 4.4.
@@
Figure 6: Comparison of our DeepDeWedge (with even/odd split) to the IsoNet, CryoCARE and the combination of IsoNet and CryoCARE for increasing noise levels on the tilt series images. We inverted the y-axes such that higher is better for all metrics.
@@
Figure 7: Slices through tomograms reconstructed with CryoCARE + IsoNet and DeepDeWedge (with even/odd split) on a tomogram from the synthetic SHREC 2021 dataset. The red lines in each slice indicate the positions of the remaining two slices. We also show the central x-z-slice through the logarithm of the magnitude of the Fourier transform of each reconstruction.
Performance Metrics.
@@
To measure the overall quality of a tomogram $\hat{\mathbf{v}}$ obtained with any of the three methods, we calculated the normalized correlation coefficient $\text{CC}(\hat{\mathbf{v}},\mathbf{v}^{*})$ between the reconstruction $\hat{\mathbf{v}}$ and the corresponding ground-truth $\mathbf{v}^{*}$, which is defined as
@@
Results.
Figure 6 shows the metrics for decreasing SNR of the tilt series. All metrics suggest that DeepDeWedge yields higher-quality reconstructions than IsoNet, CryoCARE, and CryoCARE + IsoNet.
CryoCARE achieves a lower correlation coefficient than IsoNet in the high-SNR regime, while the order is reversed for low SNR. A likely explanation is that for lower noise levels, the correlation coefficient is more sensitive to the missing wedge artifacts in the reconstructions. CryoCARE does not perform missing wedge reconstruction, so it has a lower correlation coefficient than IsoNet for higher SNR. For lower SNR, the correlation coefficient is dominated by the noise. Regarding denoising, we and others [9] have observed that CryoCARE performs better than IsoNet, which is confirmed here by looking at the correlation coefficient outside the missing wedge. As expected, CryoCARE + IsoNet combines the strengths of both methods. Compared to this combination, the overall quality of DeepDeWedge reconstructions is on par or better, depending on the noise level.
@@ 4 Discussion
In this paper, we have proposed DeepDeWedge, which is a deep-learning-based method for denoising and missing wedge reconstruction in cryo-ET.
Compared to the state-of-the-art CryoCARE + IsoNet pipeline, DeepDeWedge removes more of the noise and, therefore, produces smoother reconstructions. While this occasionally results in the loss of some high-frequency details, DeepDeWedge produces overall cleaner reconstructions with higher contrast, especially of medium or large objects such as microtubules (Figure 4, Figure 5), cellular structures (Figure 5, Figure 7), and large proteins (Figure 3, Figure 5). On the other hand, the CryoCARE + IsoNet results are more faithful to the FBP reconstructions, especially for high-frequency details, as can be seen in Figure 5. Regarding missing wedge reconstruction, which can be mainly seen in the removal of missing wedge artifacts and the reduction of artificial elongations, we find that DeepDeWedge and CryoCARE + IsoNet have overall similar performance. Both methods perform best for small and medium-sized objects like proteins or microtubules, whereas large structures like vesicles (Figure 5, Figure 7) or the boundaries of the C. reinhardtii flagella (Figure 4) pose greater challenges. This could be a result of using sub-tomograms for model fitting, as sub-tomograms may not always contain all the information that is necessary for missing wedge reconstruction of objects larger than the sub-tomogram size.
The data that DeepDeWedge fills in for the missing wedge is based on prior information about the data that the network learns during model fitting by predicting data that we artificially removed. However, it is important to keep in mind that the actual information contained in the missing wedge is irreversibly lost during tilt series acquisition.
@@
Therefore, we advise users of our method to be cautious, especially regarding reconstructing objects that are mostly perpendicular to the electron beam. Many of the Fourier components that correspond to such objects are contained in the missing wedge region. Examples of objects that are almost completely masked by the missing wedge include thin membranes or elongated proteins perpendicular to the electron beam; we discuss a concrete example and how DeepDeWedge handles it in Supplementary Information 7. It remains an open question to what extent missing-wedge-filled tomograms, whether based on deep learning or classical methods, can be trusted (see also Supplementary Information 8). To estimate the trustworthiness of a DeepDeWedge reconstruction, we recommend comparing it to a reconstruction obtained with a classical, prior-free method that is strongly data-consistent, such as FBP.
@@
Based on these considerations, we see two main applications for DeepDeWedge: First, DeepDeWedge is an effective method to improve the direct interpretability of tomograms, which may enable discoveries that are prevented by high noise levels and strong missing wedge artifacts in raw FBP reconstructions. Second, tomograms reconstructed with DeepDeWedge can be used as input for downstream tasks such as segmentation or particle picking and might improve the effectiveness of those downstream tasks. It has already been observed that denoised and/or missing-wedge-corrected tomograms can improve the performance of deep-learning-based particle pickers [11, 4, 32]. For subsequent steps, such as sub-tomogram averaging, one can then use the original raw tomograms that are not modified by the neural network.
@@
All tomograms shown in Figure 3, Figure 4, Figure 5 and Figure 7 have been deposited to Figshare and are available at https://figshare.com/articles/journal_contribution/Tomograms_shown_in_Figures_of_the_DeepDeWedge_Paper/26169538.
The EMPIAR-10045 dataset used in this study is available in the EMPIAR database under accession code 10045 (https://www.ebi.ac.uk/empiar/EMPIAR-10045/).
The tilt series collected from the C. reinhardtii flagella is available at https://download.fht.org/jug/cryoCARE/Tomo110.zip.
The EMPIAR-11078 dataset used in this study is available in the EMPIAR database under accession code 11078 (https://www.ebi.ac.uk/empiar/EMPIAR-11078/).